# OctFusion: Octree-based Diffusion Models for 3D Shape Generation

### v1

@@ 4. Experiments
We also evaluate the model generalizability of our OctFusion on Objaverse (Deitke et al., 2023), a recently proposed 3D shape dataset that contains much richer and more diverse 3D objects. For simplicity, we select a subset containing about 10k high-quality meshes in Objaverse provided by LGM(Tang et al., 2024) and train our OctFusion model with depth 10 (resolution $1024^{3}$). Fig. 9 provides unconditional generative results on Objaverse. As we can see from Fig 9, although 3D objects in Objaverse exhibit far stronger diversity than a single category of ShapeNet, our method is still capable of generating plausible 3D shapes satisfying this complex data distribution.
Figure 9. Unconditional generation results on the Objaverse dataset.
Textured Shape Generation
@@
We compare OctFusion with two recent methods: GET3D (Gao et al., 2022) and DiffTF (Cao et al., 2024). We adopt the rendering-image-based FID to evaluate the quality and diversity of the generated textured meshes. Each generated textured mesh is rendered from 20 uniformly distributed views to RGB images to compute the FID score. Table 3 and Fig. 10 provides quantitative and qualitative comparisons. Our OctFusion achieves the best FID scores on average, demonstrating the superior capability of generating high-quality and diverse textured 3D shapes. DiffTF generates NeRF as the output; we convert the generated NeRFs to 3D textured meshes using the code provided by the authors. The rendering of the extracted meshes is different from the NeRF rendering. Thus, the FID values of DiffTF are for reference.
Figure 10. The comparison of generated textured mesh of our method and other state-of-the-arts.
Table 3. The quantitative comparison of rendering-image-based FID on 3D textured mesh generation. GET3D and DiffTF were trained on 3 categories only. DiffTF is for reference as it generates NeRF instead of textured mesh.
@@
4.2. Ablations and Discussions
Octree-based Latent Representation
Here, we discuss the key benefits of our octree-based latent representation over other highly related representations, including plain octrees in LAS-Diffusion (Zheng et al., 2023) and hierarchical sparse voxels in XCube (Ren et al., 2024).
    Completeness. The leaf nodes of an octree form a complete coverage of the 3D bounding volume. We keep all octree leaf nodes in the latent space, which guarantees to contain the whole shape, whereas LAS-Diffusion and XCube prune voxels and only keep a subvolume, which may lead to holes in the generated shapes, such as the non-manifold meshes shown by the right case in Fig. 11.
    Efficiency. We observe that the node number in the octree is significantly reduced compared to LAS-Diffusion and XCube. The reason is that LAS-Diffusion needs to keep a volume shell to cover the shape surface, and XCube keeps all interior voxels of the shape. While we only keep the voxels intersecting with the surface, which makes our network more efficent than LAS-Diffusion and XCube in terms of GPU memory and inference time as shown in Table 4.
Figure 11. Some failed cases generated by LAS-Diffusion and XCube, which demonstrate the superiority of our method in ensuring the completeness and continuity.
Unified U-Net
Figure 12. Sharing weights in the U-Net for diffusion. In the 1st row are results without using shared weights, which are worse than results with sharing weights in the 2nd row.
Figure 13. The comparison of OctFusion using more diffusion stages.
Table 4. Efficiency comparison with LAS-Diffusion. The average octree node number, the GPU memory and a single forward time cost on a Nvidia 3090 GPU with batch size 1 are reported. The subscript * and ** denote first stage and second stage of corresponding methods, respectively.
@@ 4.2. Ablations and Discussions
Figure 14. The comparison of our OctFusion with AutoSDF and SDFusion on the Text2Shape dataset. Please zoom in for better inspection of the shape quality of our OctFusion and other methods.
Deeper Octree
@@
Figure 15. The sketch-conditional generation of OctFusion and LAS-Diffusion.
4.3. Applications
@@
Text-conditioned generation
We encode the textual condition with CLIP and then inject extract text features into OctFusion using cross attention. We use the Text2shape dataset (Chen et al., 2018; Cheng et al., 2023), that provides textual descriptions for the chair and table categories in ShapeNet. Fig. 14 provides qualitative comparisons between our method with AutoSDF (Mittal et al., 2022) and SDFusion (Cheng et al., 2023). It can be see that the results of AutoSDF and SDFusion exhibit severe distortions and large amounts of artifacts. On the contrary, our method is capable of generating shapes with higher quality and delicate structure details such as lots of drawers inside table while conforming to the input textual description.
Sketch-conditioned generation
We conduct sketch conditional 3D shape generation using the view-aware local attention proposed in LAS-Diffusion (Zheng et al., 2023) to aggregate sketch information to guide the generation of octrees. We train our OctFusion with the sketch dataset provided by LAS-Diffusion. Fig. 15 visualizes the results of LAS-Diffusion and our methods. Although LAS-Diffusion generates plausible results which basically conform to the geometry of input sketch, it struggles to recover the fine details of 3D shapes such as the wheel hub as well as horizontal and vertical bars of chair. On the contrary, our method is capable of possessing better geometry quality and matching better with the input sketch.
Texture Generation
@@
Figure 16. The performance of octree texture diffusion module which is capable of producing diverse texture map given a single input.
5. Conclusion

### v2

@@ 4.1.6 Model generalizability
We also evaluate the model generalizability of our OctFusion on Objaverse [Deitke2023], a recently proposed 3D shape dataset that contains much richer and more diverse 3D objects. For simplicity, we select a subset containing about 10k high-quality meshes in Objaverse provided by LGM[tang2024lgm] and train our OctFusion model with depth 10 (resolution $1024^{3}$). Fig. 8 provides unconditional generative results on Objaverse. As we can see from Fig 8, although 3D objects in Objaverse exhibit far stronger diversity than a single category of ShapeNet, our method is still capable of generating plausible 3D shapes satisfying this complex data distribution.
Figure 8: Unconditional generation results on the Objaverse dataset.
4.1.7 Textured Shape Generation
@@
We compare OctFusion with two recent methods: GET3D [Gao2022] and DiffTF [Cao2024]. We adopt the rendering-image-based FID to evaluate the quality and diversity of the generated textured meshes. Each generated textured mesh is rendered from 20 uniformly distributed views to RGB images to compute the FID score. Table 3 and Fig. 9 provides quantitative and qualitative comparisons. Our OctFusion achieves the best FID scores on average, demonstrating the superior capability of generating high-quality and diverse textured 3D shapes. DiffTF generates NeRF as the output; we convert the generated NeRFs to 3D textured meshes using the code provided by the authors. The rendering of the extracted meshes is different from the NeRF rendering. Thus, the FID values of DiffTF are for reference.
Figure 9: The comparison of generated textured mesh of our method and other state-of-the-arts.
Table 3: The quantitative comparison of rendering-image-based FID on 3D textured mesh generation. GET3D and DiffTF were trained on 3 categories only. DiffTF is for reference as it generates NeRF instead of textured mesh.
@@
Figure 10: Some failed cases generated by LAS-Diffusion and XCube, which demonstrate the superiority of our method in ensuring the completeness and continuity.
Table 4: Efficiency comparison with LAS-Diffusion and XCube. The average octree node number, the GPU memory and a single forward time cost on a Nvidia 4090 GPU with batch size 1 are reported. The subscript * and ** denote first stage and second stage of corresponding methods, respectively.
@@
Table 5: Training cost comparison with LAS-Diffusion and XCube on the model parameters, number of GPUs, and training time.
Method Params GPUs Training Time
LAS-Diffusion 57M 8 $\times$ V100 $\sim$7 days
XCube 1.6B 8 $\times$ A100 $\sim$4 days
OctFusion 33M 4 $\times$ 4090 $\sim$3 days
Figure 11: Sharing weights in the U-Net for diffusion. In the 1st row are results without using shared weights, which are worse than results with sharing weights in the 2nd row.
4.2 Ablations and Discussions
4.2.1 Octree-based Latent Representation
Here, we discuss the key benefits of our octree-based latent representation over other highly related representations, including plain octrees in LAS-Diffusion [Zheng2023] and hierarchical sparse voxels in XCube [Ren2024] and Trellis [xiang2024structured].
    Completeness. The leaf nodes of an octree form a complete coverage of the 3D bounding volume. We keep all octree leaf nodes in the latent space, which guarantees to contain the whole shape, whereas LAS-Diffusion, XCube and Trellis prune voxels and only keep a subvolume, which may lead to holes in the generated shapes, such as the non-manifold meshes shown by the right case in Fig. 10.
    Efficiency. We observe that the node number in the octree is significantly reduced compared to LAS-Diffusion and XCube. The reason is that LAS-Diffusion needs to keep a volume shell to cover the shape surface, and XCube keeps all interior voxels of the shape. While we only keep the voxels intersecting with the surface, which makes our network more efficent than LAS-Diffusion and XCube in terms of GPU memory and inference time as shown in Table 4. We also present a quantitative analysis of training costs of different methods in Table 5. OctFusion uses the lowest model parameters, GPU memory consumption and training time compared with LAS-Diffusion and XCube, achieving significant resource efficiency.
Figure 12: The comparison of our OctFusion with AutoSDF and SDFusion on the Text2Shape dataset. Please zoom in for better inspection of the shape quality of our OctFusion and other methods.
Figure 13: The comparison of OctFusion using more diffusion stages.
4.2.2 Unified U-Net
4.2.3 Deeper Octree
@@
4.3 Applications
@@
4.3.1 Text-conditioned generation
We encode the textual condition with CLIP and then inject extract text features into OctFusion using cross attention. We use the Text2shape dataset [Chen2018a, Cheng2023], that provides textual descriptions for the chair and table categories in ShapeNet. Fig. 12 provides qualitative comparisons between our method with AutoSDF [Mittal2022] and SDFusion [Cheng2023]. It can be see that the results of AutoSDF and SDFusion exhibit severe distortions and large amounts of artifacts. On the contrary, our method is capable of generating shapes with higher quality and delicate structure details such as lots of drawers inside table while conforming to the input textual description.
4.3.2 Sketch-conditioned generation
We conduct sketch conditional 3D shape generation using the view-aware local attention proposed in LAS-Diffusion [Zheng2023] to aggregate sketch information to guide the generation of octrees. We train our OctFusion with the sketch dataset provided by LAS-Diffusion. Fig. 14 visualizes the results of LAS-Diffusion and our methods. Although LAS-Diffusion generates plausible results which basically conform to the geometry of input sketch, it struggles to recover the fine details of 3D shapes such as the wheel hub as well as horizontal and vertical bars of chair. On the contrary, our method is capable of possessing better geometry quality and matching better with the input sketch.
4.3.3 Texture Generation
@@
Figure 14: The sketch-conditional generation of OctFusion and LAS-Diffusion.
Figure 15: The performance of octree texture diffusion module which is capable of producing diverse texture map given a single input.
5 Conclusion
@@
Acknowledgements
This work was supported in part by National Key R&D Program of China 2022ZD0160801, Beijing Natural Science Foundation No. 4244081, National Natural Science Foundation of China (Grant No.: 62372015), and Tencent AI Lab Rhino-Bird Focused Research Program. We also thank the anonymous reviewers for their invaluable feedback.
[…]
Appendix A Implementation Details
A.1 Textured Shape Generation
This section introduces the implementation details for textured shape generation. For data preparation, we use the same procedure in Section 4.1 to repair all the meshes in ShapeNet to obtain dense on-surface points and off-surface query points and their corresponding SDF values. Then, we reproject these points onto the surface of the original meshes and extract their corresponding RGB values to obtain colored input point clouds. We build octrees with depth 8 (resolution $256^{3}$) with colored point clouds as the input of VAE. The VAE has two separate decoders, and the latent feature dimension is set to 6 with 3 channels for geometry and another 3 channels for color. Then, we train a two-stage OctFusion. Finally, we train an additional octree-based texture diffusion model to generate color latent code based on the geometry latent code to attain a textured 3D mesh followed by the decoder of extended VAE.
A.2 Model Architecture
A.2.1 Octree-based VAE
The network architecture of the octree-based latent Variational Autoencoder (VAE), as depicted in Figure 16, is constructed via dual-octree graph convolution network. For two-stage OctFusion model, the VAE has three hierarchical levels, corresponding to octree depths of 8, 7, and 6, with corresponding resolutions of $256^{3}$, $128^{3}$, $64^{3}$. The feature dimensions are set to 24, 32 and 32 respectively.
A.2.2 OctFusion
We present the unified U-Net architecture of OctFusion in Fig. 17. The first stage, denoted as $\mathcal{F}_{1}$, is designed for generating the splitting signals, using convolutional neural network with residual connection and self-attention block. The U-Net in $\mathcal{F}_{1}$ is composed of three levels: $16^{3}$, $8^{3}$, $4^{3}$, each associated with model channels of 64, 128, and 256, respectively. For two-stage OctFusion model, the second stage $\mathcal{F}_{2}$ model is used for predicting clean latent features on each octree leaf node. The U-Net is constructed via dual octree graph convolution and has two levels $64^{3}$, $32^{3}$. The corresponding model channels are 128 and 256. At the bottom of $\mathcal{F}_{2}$ U-Net, the features are downsampled to $16^{3}$ and fed into $\mathcal{F}_{1}$.
The deeper OctFusion (such as three-stage) is shown in Fig. 18. The $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ models is used for generating splitting signals and have the same network architecture mentioned above. The $\mathcal{F}_{3}$ predicts the clean latent code on the octree generated by $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$. The U-Net of $\mathcal{F}_{3}$ has two levels $256^{3}$ and $128^{3}$, with associated model channels of 64 and 128. We also present the network architecture of octree-based texture diffusion model in Fig. 19, which is used for generating color latent codes for existing untextured 3D shape. Our octree-based texture diffusion model has the same architecture as $\mathcal{F}_{3}$ but is not unified with lower stages.
Figure 16: The network architecture of Octree-based VAE.
Figure 17: The network architecture of OctFusion U-Net.
Figure 18: The network architecture of deeper OctFusion U-Net.
Figure 19: The network architecture of OctFusion U-Net for color generation.
Appendix B Metric Definition
B.0.1 Distance
We begin by sampling points from the surfaces of both the generated mesh and the reference mesh in dataset, resulting in the point clouds denoted as $S_{g}$ and $S_{r}$, respectively. Distance between two point clouds can be evaluated by Chamfer Distance(CD) and Earth Mover’s Distance (EMD). Chamfer Distance is a symmetric measure that calculates the average distance from points in $S_{g}$ to the nearest points in $S_{r}$, and vice versa. The formula for Chamfer Distance is given by:
$\text{CD}(S_{g},S_{r})=\sum_{x\in S_{g}}\min_{y\in S_{r}}||x-y||^{2}_{2}+\sum_{y\in S_{r}}\min_{x\in S_{g}}||x-y||^{2}_{2}.$ (6)
where $d(x,y)$ is the Euclidean distance between $X$ and $Y$. Earth Mover’s Distance can be treated of as the minimum transportation from one point cloud into another. The formula of Earth Mover’s Distance (EMD) is defined as:
$\text{EMD}(S_{g},S_{r})=\min_{\phi:S_{g}\rightarrow S_{r}}\sum_{X\in S_{g}}\|X-\phi(X)\|_{2}$ (7)
where $\phi$ is a bijection.
B.0.2 Coverage (COV)
Coverage is calculated as the fraction of point clouds in the reference set that are matched to at least one point cloud in the generated set. For each point cloud in the generated set, its near neighbor in the reference set is marked as a match:
$COV(S_{g},S_{r})=\frac{|\{\text{argmin}_{Y\in S_{r}}D(X,Y)|X\in S_{g}\}|}{|S_{r}|}$ (8)
where $D(\cdot,\cdot)$ can be either CD or EMD. A high coverage score indicated that most of reference set is roughly represented within generated set.
B.0.3 Minimum Matching Distance(MMD)
Minimum Matching Distance is proposed to complement coverage as a metric that measures quality. For each point cloud in the reference set, the distance to its nearest neighbor in the generated set is computed and averaged:
$\text{MMD}(S_{g},S_{r})=\frac{1}{|S_{g}|}\sum_{Y\in S_{r}}\min_{X\in S_{g}}D(X,Y)$ (9)
where $D(\cdot,\cdot)$ can be either CD or EMD. Since MMD relies directly on the distances of the matching, it correlates well with how faithful (with respect to the reference set) the elements of generated set are.
B.0.4 1-NNA
The 1-Nearest Neighbor Assignment (1-NNA) metric evaluates the classification accuracy when employing the nearest neighbor criterion under distance measure $D$ to indicate whether a point cloud is synthetic or not. Ideally, if the generated point cloud closely mirrors the distribution of the reference set, the classification accuracy should be around 50%. The formula for 1-NNA is defined as follows.
[…]
where $N_{X}$ is the closest point cloud to $X$ under distance metric $D(\cdot,\cdot)$, and $\mathbb{1}[\cdot]$ is the indicator function.
B.0.5 shading-image-based FID
shading-image-based FID is a more robust measure for evaluating both the quality and diversity of generated shapes. To compute the FID metric, each generated shape is rendered from 20 uniformly distributed viewpoints around the shape. These rendered shading images are then used to calculate the FID scores on the rendered image set of the original training dataset. The formula for FID is defined as follows.
$\text{FID}=\frac{1}{20}\sum_{i=1}^{20}\|\mu_{g}^{i}-\mu_{r}^{i}\|^{2}+\text{Tr}(\Sigma_{g}^{i}+\Sigma_{r}^{i}-2(\Sigma_{g}^{i}\Sigma_{r}^{i})^{1/2})$ (11)
where $\mu^{i}$ and $\Sigma^{i}$ denote the mean and covariance of the $i$-th view’s shading images.
Appendix C More Results on ShapeNet
We present more unconditional generation results on ShapeNet category chair, table, airplane, car, and rifle in the following pages. These results demonstrate the quality and diversity of our proposed OctFusion.
Figure 20: More generative results on airplane
Figure 21: More generative results on car
Figure 22: More generative results on chair
Figure 23: More generative results on rifle
Figure 24: More generative results on table